# Keep It Real: Challenges in Attacking Compression-Based Adversarial Purification

### v1

@@ B.1 Adversarial Robustness
Shortly after the success of AlexNet [Krizhevsky et al., 2012], it was found that neural networks are very susceptible to adversarial attacks [Szegedy et al., 2014, Goodfellow et al., 2015]. Here, an adversary makes, usually small and imperceptible, modifications (perturbations) to, for instance, an image such that a model mislabels it.
@@
Many attacks have been developed over the years with various benefits and drawbacks. Some of the most noteworthy are FGSM (Fast Gradient Sign Method) [Goodfellow et al., 2015], iterated FGSM or iFGSM [Kurakin et al., 2017], CW (Carlini & Wagner) [Carlini and Wagner, 2017a], and PGD (Projected Gradient Decent) [Madry et al., 2019]. Additional attacks that have often been used are APGD [Croce and Hein, 2020], DeepFool [Moosavi-Dezfooli et al., 2016], and ZOO [Chen et al., 2017]. We refer the reader to the original papers for specific details, but we include the necessary details to understand this paper in Appendix˜C. The main thing to know is that PGD has a perturbation budget $\epsilon$ and number of iteration it can perform $n$. PGD then does projected gradient decent for $n$ iterations to find an adversarial example that is at most $\epsilon$ distance in $l_{\infty}$ from the original image.
@@
When the defense includes randomness, a common attack augmentation is EoT (Expectation over Transformation), in which gradients are averaged over multiple forward and backward passes [Athalye et al., 2018a, Zhang et al., 2025]. If the defense causes the gradients to be infeasible to compute, Athalye et al. [2018a] suggested Backward Pass Differentiable Approximation (BPDA), where the defense is used during the forward pass, but a differentiable approximation is used during the backward pass. The simplest is to use the identity function, while more advanced methods might train a differentiable surrogate model $g^{\prime}$ that emulates the defense $g$; i.e., $g(x)\approx g^{\prime}(x)$.
Moosavi-Dezfooli et al. [2017] show in their work “Universal adversarial perturbations” that adversarial perturbations transfer to other data samples and models; Szegedy et al. [2014] already hinted at this years before. Several later studies have explored this [Zhou et al., 2018, Zhang et al., 2021, Xia et al., 2024], attempted to explain the effect [Demontis et al., 2019], and shown the effect also exists for language models [Zou et al., 2023]. The transfer effect allows for another frequently used technique of black box attacks, where perturbations are generated for one model and applied to a black box model [Carlini et al., 2019]. Another branch of black box attacks looks only at the prediction labels of a model to generate attacks [Brendel et al., 2018].
However, as highlighted by Sheatsley et al. [2023], these are only a small set of all the attacks that could be considered. The list of possible attacks is heavily dependent on the threat model used [Carlini et al., 2019, Sheatsley et al., 2023], and if one relaxes the often-used imperceptible assumption, then there exists a wide range of semantic-preserving attacks [Chen et al., 2021, 2023, Jabary et al., 2024], and natural adversarial examples [Hendrycks and Dietterich, 2018, Hendrycks et al., 2021].
Lastly, often the best attack results are found using adaptive attacks [Carlini et al., 2019, Tramer et al., 2020, Sheatsley et al., 2023, Zhang et al., 2025]. Here, the attacks, for instance, the optimization objective, are adjusted to the defense. It is thus vital to consider adaptive attacks when a defense is resistant to standard gradient-based attacks. This work focuses on adaptive attacks with PGD as the underlying optimization method.
@@
Given the prevalence of adversarial attacks, many researchers have explored how to defend against them. We broadly view this in three groups: Architecture improvements [Singh et al., 2023, Wu et al., 2023, Fort and Lakshminarayanan, 2024], adversarial training [Szegedy et al., 2014, Goodfellow et al., 2015, Madry et al., 2019, Pang et al., 2020a, Fort and Lakshminarayanan, 2024], and adversarial purification [Nie et al., 2022, Frosio and Kautz, 2023, Zollicoffer et al., 2025, Räber et al., 2025].
The architecture branch focuses on making the models more robust by design, for instance, by taking inspiration from biology to mimic the human eye when training a ResNet model [Fort and Lakshminarayanan, 2024].
Adversarial training, as formalized by Madry et al. [2019], is perhaps the most straightforward method and consists of showing the model adversarial examples during training to ensure it classifies these correctly. This method has yielded positive results [Szegedy et al., 2014, Madry et al., 2019], but models with adversarial training can be broken [Zhang et al., 2025]. They also experience a drop in clean accuracy (accuracy on images without adversarial perturbations) [Madry et al., 2019, Carlini et al., 2019], and the robustness may not carry over to other attacks [Zhang et al., 2025].
@@
In the realm of adversarial purification, diffusion models have been proposed as a way to remove adversarial noise from input images [Nie et al., 2022, Lee and Kim, 2023]. These approaches are conceptually similar to compression-based defenses with realism: both aim to project adversarial examples back onto the manifold of natural images to restore classifier performance. However, prior work on diffusion-based purification has not explicitly investigated the role of realism as a contributing factor to robustness.
One major drawback of diffusion-based defenses is their computational cost. Diffusion models are typically an order of magnitude more expensive than VAE- or GAN-based learned compression models, making them impractical for many real-world settings [Dhariwal and Nichol, 2021, Yang and Mandt, 2023a]. This high cost limits their deployment and complicates the evaluation of robustness: mounting effective adversarial attacks against diffusion models becomes significantly harder due to the computational overhead. However, simply making gradients difficult to compute does not equate to genuine robustness. A well-designed adaptive attacker, given sufficient resources, should still be able to circumvent such defenses [Carlini et al., 2019, Lee and Kim, 2023].
Notably, the purification method proposed by Nie et al. [2022] was later defeated by Lee and Kim [2023]. However, the latter only evaluated their improved defense under the attack that broke the former. They did not develop new adaptive attacks tailored to their defense—a strategy that past research suggests would likely reduce the effectiveness of the defense. The history of adversarial robustness research shows that defenses that are not tested under strong, tailored attacks often overstate their robustness [Athalye et al., 2018a, Carlini et al., 2019, Tramer et al., 2020].
@@
The considerable interest in adversarial examples has given rise to benchmarks and leaderboards, one of which is RobustBench [Croce et al., 2021]. This provides a leaderboard of models that are supposed to be robust to adversarial examples. However, note that Fort and Lakshminarayanan [2024] claimed their model beat the RobustBench leaderboard, but, using adaptive attacks, Zhang et al. [2025] showed that the model is still very vulnerable to adversarial examples.
We will use RobustBench as a source of robust models, and thus enable us to test if compression-based purification would help as claimed in [Räber et al., 2025]. For this, we take top-performing models from the leaderboard with open-sourced model weights and test whether the compression-based purification defenses improve the models’ robustness.
@@ B.2 Realism
However, metrics such as PSNR, MS-SSIM [Wang et al., 2003], or LPIPS [Zhang et al., 2018] correlate poorly with human perception, and directly quantifying perceptual distortion remains a challenging problem. Instead of a perceptual distortion metric, one can also measure both distortion and realism and optimize the compression model for both. Realism can be formally defined as:
@@
where $d(\cdot,\cdot)$ is a divergence measure, such as Kullback-Leibler. Unlike distortion, realism is a no-reference metric, requiring only the generated image distribution to match that of natural images. Although measuring realism remains challenging [Theis, 2024], a widely used proxy is the Fréchet Inception Distance (FID) [Heusel et al., 2017], which compares the distributions extracted by a pretrained Inception network. FID has gained popularity for capturing both fidelity and diversity in generated samples.
@@ B.3 Compression Models
End-to-End Optimized Image Compression [Ballé et al., 2016] jointly learns analysis and synthesis transforms along with an entropy model over quantized latents, directly optimizing rate–distortion performance in a VAE-style framework. The Hyperprior model [Ballé et al., 2018] extends this approach by introducing a secondary network that predicts spatially adaptive Gaussian scales, better capturing local image statistics. HiFiC [Mentzer et al., 2020] enhances learned compression with GANs to produce realistic reconstructions. By refining network design and training with perceptual losses, it generates outputs that closely resemble natural images.
ELIC [He et al., 2022a] introduced uneven channel grouping and deeper autoregressive context networks for more accurate entropy modeling, yielding faster convergence and lower bitrates. PO-ELIC [He et al., 2022b] then augmented this foundation with adversarial fine-tuning and perceptual losses to enrich texture realism at low bitrates.
More recent work has shifted toward models that offer explicit, user-controllable trade-offs between rate, distortion, and realism within a single network. MRIC [Agustsson et al., 2023] introduced a conditional generator that, at a fixed bit rate, lets users interpolate between low-distortion and high-realism reconstructions by tuning a realism flag $\beta$, thereby explicitly navigating the distortion–realism trade-off within a single model. CRDR [Iwai et al., 2024] built on this by adding a discrete quality-level input $q$ alongside $\beta$, and embedding interpolation channel attention layers in both encoder and generator to yield true variable-rate compression, allowing joint control over bitrate, distortion, and realism with one network.
Diffusion models are great at generating images, which is also taken advantage of in image compression. Models [Yang and Mandt, 2023b, Careil et al., 2023, Hoogeboom et al., 2023, Xu et al., 2024, Relic et al., 2024] that focus on extra low bitrate or extra high realism without worrying about compute use diffusion. These methods are orders of magnitude more expensive than VAE-based methods, but achieve great realism.
Reducing computational complexity has also been a focus in compression with the line of work like Cool-Chic [Ladune et al., 2023] and C3 [Kim et al., 2024]. These methods overfit a latent, auto-regressive latent model and decoder to a single image, and send all three over the channel. High realism can be achieved at this level of complexity [Ballé et al., 2024], but as these models are INRs, computing a gradient through the compression model is not straightforward, and encoding is computationally expensive.
@@ B.4 Compression as adversarial defense
Using compression as a defense for neural networks is not a new idea. It has been explored for almost a decade [Guo et al., 2017, Dziugaite et al., 2016, Das et al., 2017, Jia et al., 2019, Ferrari et al., 2023] where some authors also explored using iterated compression and decompression cycles [Ferrari et al., 2023, Räber et al., 2025]. And more recently, using video compression to defend video classifiers [Song et al., 2024] and model quantization to defend general models [Costa and Pinto, 2024].
@@ Appendix C Details on Adversarial Attacks
PGD is a very strong attack, and most defenses in a white-box setting can be broken by it when using the right objective. [^1] Personal communication with Nicholas Carlini with further evidence in [Zhang et al., 2025].
@@ Appendix F Training U-Nets for BPDA
The U-Net used to approximate the compression and decompression step in the BPDA attack follows a standard U-Net architecture [Ronneberger et al., 2015]. We use two downsampling layers, which transform the input images from a resolution of (3,224,224) to an intermediate representation of (256,56,56). For each experiment, we train the U-Net on the whole dataset used in the experiment. The U-Net is trained for 20 epochs using the L1-loss between the input image (original image with or without adversarial noise) and the target image (reconstruction of the input image) using an Adam optimizer with a learning rate of 0.001 and a learning rate schedule. The learning rate scheduler is StepLR from PyTorch, configured with a step size of 5 and a gamma of 0.1. During an epoch, each batch is used 4 times, 3 passes with additional noise added to the images to simulate adversarial noise, and one clean pass with the original images.
@@ G.1 Attacking Adversarial Purification
We ran a limited attack on a diffusion-based adversarial purification defense [Lee and Kim, 2023]. Since running (and attacking) this defense is more computationally expensive compared to the compression models used in the other results shown in this paper, we focused on a smaller part of the ImageNet dataset.
@@
Table 10: Accuracy of the U-net attack on adversarial purification. The PGD + EOT are from [Lee and Kim, 2023].
We ran a U-Net BPDA attack on 100 images. Table˜10 shows the accuracy of an attack with epsilon 8/255. The attacked accuracy of 69% shows that the U-Net attack was not able to produce practical gradients, as the accuracy is higher than the base accuracy. The attack proposed by [Lee and Kim, 2023] showed an accuracy of 42.15% with an even lower epsilon of 4/255. In addition, we apply a very aggressive attack with an epsilon of 64/255. However, Table˜10 shows that the results are comparable to adding random noise to the image, showing that the gradients carry no usable signal.
@@
We also ran the PGD + EOT attack from [Lee and Kim, 2023] for epsilon 8/255 and 100 PGD iterations. The weaker defnse we attacked used just 9 diffusion steps. This allows us to compute the gradient through the entire defense. To better evaluate our results, we compute the accuracy of the full defense + classifier after every PGD iteration. We started the attack with 64 correctly classified images. Figure˜5 shows the decrease in accuracy over 100 PGD iterations. The non-deterministic nature of the diffusion model leads to a certain variance in the accuracy at higher iterations. We therefore report the average accuracy of 28.1% over iterations 50 to 100 and the minimal accuracy of 21.9%. The bottom graph in Figure˜5 shows the percentage of images that have never been missclassified up to the current PGD iteration. Evaluating this metric provides an estimate of worst-case adversarial robustness, which is 14.1% after 100 iterations.
@@ G.3 RobustBench
Instead of evaluating against standard (adversarially weak) classifiers, one can use pretrained robust classifiers to assess whether realism offers additional benefits in already robust settings. We ran the compression-based defense on the top 11 models from RobustBench [Singh et al., 2023, Bai et al., 2024, Amini et al., 2024, Xu et al., 2025, Liu et al., 2025]. The results in Table˜11 show that applying CRDR with high quality and high realism does not improve performance—in fact, it consistently reduces accuracy across all tested robust models. While CRDR benefits standard models by projecting inputs back onto the natural image manifold, we see two reasons for these results. 1) The information loss and subtle degradations introduced by compression can harm standard and robust accuracy when applied to already robust models, as the compression may discard the specific features these robust models have learned to utilize for classification, explaining why performance decreases rather than increases. 2) The robustified models have been overfitted to ImageNet images ([Torralba and Efros, 2011, Liu and He, 2024]) without any noise and images with exactly the type of noise PGD produces. We leave it for future work to resolve which, if any, of these explanations are correct.
@@
Amini ConvNeXt-L [Amini et al., 2024] 78.58 76.04 61.98 57.08 43.98 36.20 18.30 11.84
Amini Swin-L [Amini et al., 2024] 78.98 77.10 65.12 60.18 49.42 41.66 25.06 17.28
Bai NUTS [Bai et al., 2024] 81.48 80.12 70.66 66.76 51.10 46.04 14.28 11.62
Liu ConvNeXt-B [Liu et al., 2025] 77.16 74.48 58.40 53.34 39.44 33.06 18.34 13.48
Liu ConvNeXt-L [Liu et al., 2025] 78.62 76.18 60.48 55.26 41.44 34.32 19.80 14.24
Liu Swin-B [Liu et al., 2025] 76.78 74.62 59.80 53.80 41.08 34.60 19.80 13.84
Liu Swin-L [Liu et al., 2025] 79.00 77.12 61.94 57.04 43.04 36.76 22.24 16.00
Singh ConvNeXt-B [Singh et al., 2023] 75.94 73.52 58.00 52.18 38.84 32.62 16.18 12.08
Singh ConvNeXt-L [Singh et al., 2023] 77.66 75.16 60.32 54.84 41.86 35.24 19.12 15.02
Xu Swin-B [Xu et al., 2025] 77.26 74.90 58.58 53.62 39.18 32.88 16.32 12.40
Xu Swin-L [Xu et al., 2025] 79.40 77.30 62.32 57.00 42.20 35.86 19.06 14.72

### v3

@@ B.1 Adversarial Robustness
Shortly after the success of AlexNet [8], it was found that neural networks are very susceptible to adversarial attacks [9, 10]. Here, an adversary makes, usually small and imperceptible, modifications (perturbations) to, for instance, an image such that a model mislabels it.
@@
Many attacks have been developed over the years with various benefits and drawbacks. Some of the most noteworthy are FGSM (Fast Gradient Sign Method) [10], iterated FGSM or iFGSM [11], CW (Carlini & Wagner) [12], and PGD (Projected Gradient Decent) [13]. Additional attacks that have often been used are APGD [59], DeepFool [60], and ZOO [61]. We refer the reader to the original papers for specific details, but we include the necessary details to understand this paper in Appendix˜C. The main thing to know is that PGD has a perturbation budget $\epsilon$ and number of iteration it can perform $n$. PGD then does projected gradient decent for $n$ iterations to find an adversarial example that is at most $\epsilon$ distance in $l_{\infty}$ from the original image.
@@
When the defense includes randomness, a common attack augmentation is EoT (Expectation over Transformation), in which gradients are averaged over multiple forward and backward passes [14, 7]. If the defense causes the gradients to be infeasible to compute, Athalye et al. [14] suggested Backward Pass Differentiable Approximation (BPDA), where the defense is used during the forward pass, but a differentiable approximation is used during the backward pass. The simplest is to use the identity function, while more advanced methods might train a differentiable surrogate model $g^{\prime}$ that emulates the defense $g$; i.e., $g(x)\approx g^{\prime}(x)$.
Moosavi-Dezfooli et al. [62] show in their work “Universal adversarial perturbations” that adversarial perturbations transfer to other data samples and models; Szegedy et al. [9] already hinted at this years before. Several later studies have explored this [63, 64, 65], attempted to explain the effect [66], and shown the effect also exists for language models [67]. The transfer effect allows for another frequently used technique of black box attacks, where perturbations are generated for one model and applied to a black box model [4]. Another branch of black box attacks looks only at the prediction labels of a model to generate attacks [68].
However, as highlighted by Sheatsley et al. [6], these are only a small set of all the attacks that could be considered. The list of possible attacks is heavily dependent on the threat model used [4, 6], and if one relaxes the often-used imperceptible assumption, then there exists a wide range of semantic-preserving attacks [69, 70, 71], and natural adversarial examples [72, 73].
Lastly, often the best attack results are found using adaptive attacks [4, 5, 6, 7]. Here, the attacks, for instance, the optimization objective, are adjusted to the defense. It is thus vital to consider adaptive attacks when a defense is resistant to standard gradient-based attacks. This work focuses on adaptive attacks with PGD as the underlying optimization method.
@@
Given the prevalence of adversarial attacks, many researchers have explored how to defend against them. We broadly view this in three groups: Architecture improvements [15, 16, 17], adversarial training [9, 10, 13, 18, 17], and adversarial purification [19, 20, 21, 2].
The architecture branch focuses on making the models more robust by design, for instance, by taking inspiration from biology to mimic the human eye when training a ResNet model [17].
Adversarial training, as formalized by Madry et al. [13], is perhaps the most straightforward method and consists of showing the model adversarial examples during training to ensure it classifies these correctly. This method has yielded positive results [9, 13], but models with adversarial training can be broken [7]. They also experience a drop in clean accuracy (accuracy on images without adversarial perturbations) [13, 4], and the robustness may not carry over to other attacks [7].
@@
In the realm of adversarial purification, diffusion models have been proposed as a way to remove adversarial noise from input images [19, 23]. These approaches are conceptually similar to compression-based defenses with realism: both aim to project adversarial examples back onto the manifold of natural images to restore classifier performance. However, prior work on diffusion-based purification has not explicitly investigated the role of realism as a contributing factor to robustness.
One major drawback of diffusion-based defenses is their computational cost. Diffusion models are typically an order of magnitude more expensive than VAE- or GAN-based learned compression models, making them impractical for many real-world settings [24, 25]. This high cost limits their deployment and complicates the evaluation of robustness: mounting effective adversarial attacks against diffusion models becomes significantly harder due to the computational overhead. However, simply making gradients difficult to compute does not equate to genuine robustness. A well-designed adaptive attacker, given sufficient resources, should still be able to circumvent such defenses [4, 23].
Notably, the purification method proposed by Nie et al. [19] was later defeated by Lee and Kim [23]. However, the latter only evaluated their improved defense under the attack that broke the former. They did not develop new adaptive attacks tailored to their defense—a strategy that past research suggests would likely reduce the effectiveness of the defense. The history of adversarial robustness research shows that defenses that are not tested under strong, tailored attacks often overstate their robustness [14, 4, 5].
@@
The considerable interest in adversarial examples has given rise to benchmarks and leaderboards, one of which is RobustBench [54]. This provides a leaderboard of models that are supposed to be robust to adversarial examples. However, note that Fort and Lakshminarayanan [17] claimed their model beat the RobustBench leaderboard, but, using adaptive attacks, Zhang et al. [7] showed that the model is still very vulnerable to adversarial examples.
We will use RobustBench as a source of robust models, and thus enable us to test if compression-based purification would help as claimed in [2]. For this, we take top-performing models from the leaderboard with open-sourced model weights and test whether the compression-based purification defenses improve the models’ robustness.
@@ B.2 Realism
However, metrics such as PSNR, MS-SSIM [26], or LPIPS [27] correlate poorly with human perception, and directly quantifying perceptual distortion remains a challenging problem. Instead of a perceptual distortion metric, one can also measure both distortion and realism and optimize the compression model for both. Realism can be formally defined as:
@@
where $d(\cdot,\cdot)$ is a divergence measure, such as Kullback-Leibler. Unlike distortion, realism is a no-reference metric, requiring only the generated image distribution to match that of natural images. Although measuring realism remains challenging [28], a widely used proxy is the Fréchet Inception Distance (FID) [29], which compares the distributions extracted by a pretrained Inception network. FID has gained popularity for capturing both fidelity and diversity in generated samples.
@@ B.3 Compression Models
End-to-End Optimized Image Compression [31] jointly learns analysis and synthesis transforms along with an entropy model over quantized latents, directly optimizing rate–distortion performance in a VAE-style framework. The Hyperprior model [45] extends this approach by introducing a secondary network that predicts spatially adaptive Gaussian scales, better capturing local image statistics. HiFiC [33] enhances learned compression with GANs to produce realistic reconstructions. By refining network design and training with perceptual losses, it generates outputs that closely resemble natural images.
ELIC [32] introduced uneven channel grouping and deeper autoregressive context networks for more accurate entropy modeling, yielding faster convergence and lower bitrates. PO‑ELIC [34] then augmented this foundation with adversarial fine‑tuning and perceptual losses to enrich texture realism at low bitrates.
More recent work has shifted toward models that offer explicit, user‑controllable trade‑offs between rate, distortion, and realism within a single network. MRIC [37] introduced a conditional generator that, at a fixed bit rate, lets users interpolate between low‑distortion and high‑realism reconstructions by tuning a realism flag $\beta$, thereby explicitly navigating the distortion–realism trade‑off within a single model. CRDR [38] built on this by adding a discrete quality‑level input $q$ alongside $\beta$, and embedding interpolation channel attention layers in both encoder and generator to yield true variable‑rate compression, allowing joint control over bitrate, distortion, and realism with one network.
Diffusion models are great at generating images, which is also taken advantage of in image compression. Models [75, 76, 77, 78, 79] that focus on extra low bitrate or extra high realism without worrying about compute use diffusion. These methods are orders of magnitude more expensive than VAE-based methods, but achieve great realism.
Reducing computational complexity has also been a focus in compression with the line of work like Cool-Chic [80] and C3 [81]. These methods overfit a latent, auto-regressive latent model and decoder to a single image, and send all three over the channel. High realism can be achieved at this level of complexity [82], but as these models are INRs, computing a gradient through the compression model is not straightforward, and encoding is computationally expensive.
@@ B.4 Compression as adversarial defense
Using compression as a defense for neural networks is not a new idea. It has been explored for almost a decade [1, 39, 40, 41, 42] where some authors also explored using iterated compression and decompression cycles [42, 2]. And more recently, using video compression to defend video classifiers [83] and model quantization to defend general models [84].
@@ Appendix C Details on Adversarial Attacks
PGD is a very strong attack, and most defenses in a white-box setting can be broken by it when using the right objective. [^1] Personal communication with Nicholas Carlini with further evidence in [7].
@@ Appendix F Training U-Nets for BPDA
The U-Net used to approximate the compression and decompression step in the BPDA attack follows a standard U-Net architecture [50]. We use two downsampling layers, which transform the input images from a resolution of (3,224,224) to an intermediate representation of (256,56,56). For each experiment, we train the U-Net on the whole dataset used in the experiment. The U-Net is trained for 20 epochs using the L1-loss between the input image (original image with or without adversarial noise) and the target image (reconstruction of the input image) using an Adam optimizer with a learning rate of 0.001 and a learning rate schedule. The learning rate scheduler is StepLR from PyTorch, configured with a step size of 5 and a gamma of 0.1. During an epoch, each batch is used 4 times, 3 passes with additional noise added to the images to simulate adversarial noise, and one clean pass with the original images.
@@ G.1 Attacking Adversarial Purification
We ran a limited attack on a diffusion-based adversarial purification defense [23]. Since running (and attacking) this defense is more computationally expensive compared to the compression models used in the other results shown in this paper, we focused on a smaller part of the ImageNet dataset.
@@
Table 10: Accuracy of the U-net attack on adversarial purification. The PGD + EOT are from [23].
We ran a U-Net BPDA attack on 100 images. Table˜10 shows the accuracy of an attack with epsilon 8/255. The attacked accuracy of 69% shows that the U-Net attack was not able to produce practical gradients, as the accuracy is higher than the base accuracy. The attack proposed by [23] showed an accuracy of 42.15% with an even lower epsilon of 4/255. In addition, we apply a very aggressive attack with an epsilon of 64/255. However, Table˜10 shows that the results are comparable to adding random noise to the image, showing that the gradients carry no usable signal.
@@
We also ran the PGD + EOT attack from [23] for epsilon 8/255 and 100 PGD iterations. The weaker defnse we attacked used just 9 diffusion steps. This allows us to compute the gradient through the entire defense. To better evaluate our results, we compute the accuracy of the full defense + classifier after every PGD iteration. We started the attack with 64 correctly classified images. Figure˜5 shows the decrease in accuracy over 100 PGD iterations. The non-deterministic nature of the diffusion model leads to a certain variance in the accuracy at higher iterations. We therefore report the average accuracy of 28.1% over iterations 50 to 100 and the minimal accuracy of 21.9%. The bottom graph in Figure˜5 shows the percentage of images that have never been missclassified up to the current PGD iteration. Evaluating this metric provides an estimate of worst-case adversarial robustness, which is 14.1% after 100 iterations.
@@ G.3 RobustBench
Instead of evaluating against standard (adversarially weak) classifiers, one can use pretrained robust classifiers to assess whether realism offers additional benefits in already robust settings. We ran the compression-based defense on the top 11 models from RobustBench [15, 55, 56, 57, 58]. The results in Table˜11 show that applying CRDR with high quality and high realism does not improve performance—in fact, it consistently reduces accuracy across all tested robust models. While CRDR benefits standard models by projecting inputs back onto the natural image manifold, we see two reasons for these results. 1) The information loss and subtle degradations introduced by compression can harm standard and robust accuracy when applied to already robust models, as the compression may discard the specific features these robust models have learned to utilize for classification, explaining why performance decreases rather than increases. 2) The robustified models have been overfitted to ImageNet images ([86, 87]) without any noise and images with exactly the type of noise PGD produces. We leave it for future work to resolve which, if any, of these explanations are correct.
@@
Amini ConvNeXt-L [56] 78.58 76.04 61.98 57.08 43.98 36.20 18.30 11.84
Amini Swin-L [56] 78.98 77.10 65.12 60.18 49.42 41.66 25.06 17.28
Bai NUTS [55] 81.48 80.12 70.66 66.76 51.10 46.04 14.28 11.62
Liu ConvNeXt-B [58] 77.16 74.48 58.40 53.34 39.44 33.06 18.34 13.48
Liu ConvNeXt-L [58] 78.62 76.18 60.48 55.26 41.44 34.32 19.80 14.24
Liu Swin-B [58] 76.78 74.62 59.80 53.80 41.08 34.60 19.80 13.84
Liu Swin-L [58] 79.00 77.12 61.94 57.04 43.04 36.76 22.24 16.00
Singh ConvNeXt-B [15] 75.94 73.52 58.00 52.18 38.84 32.62 16.18 12.08
Singh ConvNeXt-L [15] 77.66 75.16 60.32 54.84 41.86 35.24 19.12 15.02
Xu Swin-B [57] 77.26 74.90 58.58 53.62 39.18 32.88 16.32 12.40
Xu Swin-L [57] 79.40 77.30 62.32 57.00 42.20 35.86 19.06 14.72